The Creative Publishing Company (CPC) is a coupon book publisher with markets in several
southeastern states. CPC coupon books are either sold directly to the public, sold through
religious and other charitable organizations, or given away as promotional items. Operating
experience during the past year suggests the following demand function for CPC’s coupon
books:
Q = 5,000 – 4,000P + 0.02Pop + 0.5I + 1.5A
Where: Q is quantity, P is price ($), Pop is population, I is disposable income per household ($),
and A is advertising expenditures ($).
A. Determine the demand faced by CPC in a typical market in which P = $10, Pop = 1,000,000
persons, I = $30,000, and A = $10,000.
B. Calculate the level of demand if CPC increases annual advertising expenditures from
$10,000 to $15,000.
C. Calculate the demand curves faced by CPC in parts A and B.
1
Expert's answer
2017-10-09T15:08:07-0400
A. Q = 5,000 – 4,000P + 0.02Pop + 0.5I + 1.5A
So, quantity of demand is:
Q = 5,000 – 4,000*10 + 0.02*1,000,000 + 0.5*30,000+1.5*10,000 = 15,000 (coupon books)
B. Q = 5,000 – 4,000*10 + 0.02*1,000,000 + 0.5*30,000+1.5*15,000 = 22,500 (coupon books)
C. Demand curves are:
Q1 = 5,000 – 4,000*P + 0.02*1,000,000 + 0.5*30,000+1.5*10,000 = 190,000 – 4,000P
Q2 = 5,000 – 4,000*P + 0.02*1,000,000 + 0.5*30,000+1.5*15,000 = 197,500 – 4,000P